All Geometries of the Mathieu Group M_{11} Based on Maximal Subgroups

Dimitri Leemans

Abstract

Using a Cayley program, we get all firm, residually connected geometries whose rank-two residues satisfy the intersection property, on which M_{11} acts flag-transitively, and in which the stabilizer of each element is a maximal subgroup of M_{11}.